### v2

# Who Should I Listen To? Adaptive Collaboration in Personalized Federated Learning

Abstract
Data heterogeneity is a central challenge in federated learning, and personalized federated learning (PFL) aims to address it by tailoring models to each client’s distribution. Yet many PFL methods fail to consistently outperform local or centralized baselines, suggesting a mismatch between the collaboration they enforce and the structure of the data. We propose an approach based on adaptive collaboration, where clients decide adaptively not only how much to rely on others, but also whom to trust—at the level of individual examples. We instantiate this principle in FedMosaic, a federated co-training method in which clients exchange predictions over a shared unlabeled dataset. This enables fine-grained trust decisions that are difficult to achieve with parameter sharing alone. Each client adjusts its loss weighting based on the agreement between private and public data, and contributes to global pseudo-labels in proportion to its estimated per-example confidence. Empirically, FedMosaic improves upon state-of-the-art PFL methods across diverse non-IID settings, and we provide convergence guarantees under standard assumptions. Our results demonstrate the potential of data-aware collaboration for robust and effective personalization.
1 Introduction
Table 1: Average test Accuracy on DomainNet and Office dataset (details in sec.4). Most personalized FL methods fail to surpass local training baseline. FedMosaic consistently exceeds both core baselines through adaptive, example-level collaboration. Color Map: baselines, worse than baselines, worse than local training, better than baselines.
Method DomainNet Office
Centralized $0.6624$ $(0.004)$ $0.4092$ $(0.006)$
FL FedAvg $0.3100$ $(0.008)$ $0.3725$ $(0.008)$
FedProx $0.5523$ $(0.001)$ $0.5839$ $(0.003)$
PFL Per-FedAvg $0.7248$ $(0.004)$ $0.7192$ $(0.005)$
pFedMe $0.7521$ $(0.005)$ $0.7483$ $(0.007)$
APFL $0.8059$ $(0.003)$ $0.8091$ $(0.001)$
FedPHP $0.7825$ $(0.006)$ $0.7636$ $(0.004)$
Local Training $0.8464$ $(0.001)$ $0.8679$ $(0.004)$
FedMosaic $\mathbf{0.8744}$ $(0.0002)$ $\mathbf{0.8906}$ $(0.0001)$
Federated learning promises collaboration without compromising privacy, but when client data is heterogeneous, standard approaches often collapse. Personalized federated learning (PFL) seeks to address this challenge, yet much of the field rests on shaky foundations. While the task of PFL is well-defined in principle, i.e., optimizing performance on each client’s distribution, many existing methods are evaluated under ill-posed experimental conditions. Several fail to compare against purely local training (Fallah et al., 2020; T Dinh et al., 2020; Deng et al., 2020; Li et al., 2021d), where clients learn independently without communication. If a PFL method underperforms this baseline, collaboration is not only unnecessary but actively detrimental (cf. Tab. 1). Similarly, many approaches omit comparisons to centralized training, which, while often infeasible due to privacy constraints, reveals whether a global model could succeed despite heterogeneity. In cases where centralized training performs well, the real problem lies not in the need for personalization, but in the failure of existing federated aggregation schemes under distribution shift.
In personalized federated learning, each client $i\in[N]$ has access to a dataset $\mathcal{D}_{i}$ drawn from a local data distribution $\mathcal{D}_{i}$, which may vary across clients. The goal is to learn personalized models $h_{1},\dots,h_{N}$ that minimize the local risk at each client, that is,
$\min_{h_{1},\dots,h_{N}}\sum_{i=1}^{N}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\left[\mathcal{L}\left(h(x),y\right)\right]\enspace.$
This objective highlights a subtle but crucial point: since clients are evaluated on their own distributions, local training can sometimes outperform centralized training. Thus, strong local baselines must be considered when assessing the true benefit of personalization.
One might object that centralized models optimize global performance, while PFL focuses on client-specific distributions. However, this overlooks a key tradeoff. Centralized models benefit from far more data, which can reduce variance and improve generalization, even on heterogeneous test sets. Personalized models may better align with individual distributions, but often suffer from data scarcity and overfitting. In practice, centralized training often outperforms PFL methods when the gains from data volume outweigh the benefits of specialization. For this reason, centralized training remains a critical reference point, even when personalization is the objective.
This confusion arises because personalization is often proposed as a vague remedy for heterogeneity without clarifying when it is needed, how it should work, or what the right metrics for success are. We argue that progress in PFL requires a clearer conceptual foundation. Personalization should not be the default; it should emerge from an understanding of what each client needs and how collaboration benefits them. In particular, a meaningful PFL solution must adapt the degree of collaboration based on the client’s data and context, rather than enforcing a uniform global training scheme. It must also account for heterogeneity not only at the client level but at the level of individual examples or regions of the data space, since clients may align on some concepts and diverge on others. Without such granularity, personalization reduces to blunt model-level adjustments that fail to capture the true complexity of data heterogeneity.
These considerations point to a general principle: effective personalization in federated learning must enable adaptive, data-specific collaboration. A PFL method should allow clients to modulate how much they learn from others based on their own needs, while supporting fine-grained interactions that reflect the structure of the data itself. While this principle could be realized in many ways, one particularly promising direction comes from federated semi-supervised learning, e.g., distributed distillation (Bistritz et al., 2020) or federated co-training (Abourayya et al., 2025). Unlike classical methods that share model parameters, these methods enable clients to interact by sharing predictions on a shared, unlabeled public dataset. This shift makes it possible to personalize collaboration per example: a client can contribute only to those examples for which it has expertise, and consensus mechanisms can weight contributions accordingly. Such mechanisms naturally reflect the varying reliability of clients across the data space. Two clients who have seen many cat images might agree on how to label a new cat photo, while a client that has never seen dogs should not influence the labeling of dog images. This level of collaboration is fundamentally difficult—if not impossible—to achieve through parameter aggregation alone.
In this work, we demonstrate that this principle can be put into practice. We propose a method called FedMosaic, a personalized federated co-training algorithm that enables adaptive, fine-grained collaboration through prediction sharing. FedMosaic uses two key mechanisms: a dynamic weighting strategy that allows each client to adjust its reliance on global versus local signals, and an expertise-aware consensus mechanism that weights client contributions according to their demonstrated competence on different parts of the data. Both mechanisms operate on predictions over a public dataset, enabling personalization that is responsive to real distributional structure rather than imposed heuristically.
While FedMosaic achieves strong empirical performance, outperforming state-of-the-art personalized FL methods as well as both local and centralized baselines, its core contribution is conceptual. It shows that meaningful personalization is achievable when we rethink the way clients interact. Rather than pushing more complex model-sharing schemes or training tricks, we argue for a shift in perspective: personalization is not a layer added on top of existing FL algorithms, but a question of how collaboration happens. FedMosaic is one instantiation of this broader view, illustrating that principled, adaptive collaboration can substantially advance personalized federated learning. We hope this perspective inspires future research on data-driven personalization and more nuanced models of federated interaction.
2 Related Work
Federated Learning (FL) aims to train models collaboratively across decentralized clients without compromising data privacy. However, heterogeneous data distributions across clients (non-IID settings) present a persistent challenge that degrades performance. Approaches addressing heterogeneity broadly fall into two categories: traditional FL and personalized FL (PFL) methods. We review these groups in relation to our method, FEDMOSAIC.
Traditional Federated Learning:
Traditional federated learning methods typically learn a single global model. FedAvg (McMahan et al., 2017) averages local models but struggles under non-IID data due to client drift. Subsequent methods attempt to correct this: SCAFFOLD (Karimireddy et al., 2020) uses control variates to correct the local updates, FedProx (Li et al., 2020) adds a proximal term to each client’s loss function to stabilize training, and FedDyn (Acar et al., 2021) introduces dynamic regularization. Others use representation alignment, such as MOON (Li et al., 2021a), which applies a contrastive loss to align local and global features. These methods implicitly assume a global model can suffice, which may fail under strong heterogeneity. Moreover, parameter sharing can pose privacy risks (Zhu et al., 2019; Abourayya et al., 2025).
Personalized Federated learning (PFL)
Personalized Federated learning methods tailor models to individual clients, addressing non-IID challenges through different strategies.
Meta-learning and Regularization-Based Methods: The methods optimize a shared initialization or constrain local updates. For example, Per-FedAvg (Fallah et al., 2020)uses MAML to learn a shared initialization, while Ditto (Li et al., 2021b) regularizes local updates toward a global reference. PFedMe(T Dinh et al., 2020) applies bi-level optimization to decouple personalization from global learning.
Personalized Aggregation strategies: These methods dynamically aggregate models based on client similarity or adaptive weighting. APFL (Deng et al., 2020) introduces an adaptive mixture of global and local models, allowing clients to interpolate between shared and personalized parameters based on their data distribution. FedAMP(Huang et al., 2021) uses attention to weight client contributions based on similarity. Other methods select collaborators (e.g., FedFomo (Zhang et al., ), FedPHP (Li et al., 2021d)) or apply layer-wise attention (FedALA (Zhang et al., 2023c)).
Model Splitting Architectures: These approaches partition models into shared and personalized components. FedPer (Arivazhagan et al., 2019)eeps shared base layers and personalizes top layers. FedRep shares a backbone but personalizes the head.(Collins et al., 2021) shares a backbone but personalizes the head. FedBN (Li et al., 2021c) personalizes batch normalization layers to tackle feature shift. Other recent methods such as FedAS (Yang et al., 2024), GPFL (Zhang et al., 2023b), and FedBABU (Oh et al., 2021) disentangle or freeze specific parts of the model to balance generalization and personalization. PFedHN (Shamsian et al., 2021) uses a hypernetwork that generates personalized model parameters conditioned on client identity.
Knowledge Distillation Approaches: Knowledge distillation methods transfer knowledge from global or peer models to personalized local models. FedProto (Tan et al., 2022) aligns class-wise feature prototypes across clients, FedPAC (Xu et al., 2023) uses contrastive learning to distill knowledge into personalized models, and FedKD (Wu et al., 2022) reduces communication cost by distilling knowledge from a teacher ensemble to lightweight client models. FedMatch (Chen et al., 2021) uses consistency regularization to unlabeled and noisy data, FedDF (Lin et al., 2020) aggregates predictions via ensemble distillation, and FedNoisy (Liang et al., 2023) focuses on robust aggregation in the presence of noisy labels or adversarial participants.
Despite this progress, existing PFL methods often share several limitations: (i) Static collaboration: Most PFL methods rely on fixed rules (e.g., aggregation weights or model splits), lacking adaptivity to client-specific or example-level variation. (ii) Privacy risks: Sharing model parameters, gradients, or even soft labels may expose sensitive information. (iii) Limited generality: Many methods are tailored to specific heterogeneity types (e.g., label skew in case of FedMix, or feature shift in case of FedBN). (iv) Communication/computational overhead: Some require complex multi-model training or costly synchronization. To overcome these limitations, we argue that PFL methods should use some form of dynamic modulation and per-example trust weighting. We propose a method that achieves both while sharing only hard predictions on an unlabeled dataset to enhance privacy and minimize overhead.
3 Personalized Federated Co-Training: Adaptive and Expert-Aware Collaboration
We now introduce FedMosaic as a concrete realization of the principle proposed in Sec. 1, i.e., that effective personalization in federated learning arises from adaptive, data-specific collaboration.
It builds on the idea of federated co-training, a privacy-friendly alternative to parameter sharing. In federated co-training (Abourayya et al., 2025), each client shares only its predictions (hard labels) on a shared, unlabeled public dataset. These predictions are aggregated into a consensus pseudo-labeling, which is then used to augment the local datasets. This approach has been shown to substantially improve privacy. Therefore, in our setting we assume a shared unlabeled dataset $U\subset\mathcal{X}\times\mathcal{Y}$. Such datasets are often available or can be synthesized, for example from public data repositories or via generative models. Clients share only their predictions on $U$ and no sensitive information or parameters are exchanged.
However, this basic setup introduces two key challenges. (i) When to trust the global signal? Clients may disagree with the global consensus, particularly when their data distribution diverges from the majority. Blindly following the global pseudo-labels can harm performance. (ii) Whose predictions to trust? Not all clients are equally competent across all regions of the data space. A good method should rely more on clients that are confident and accurate for specific examples.
To address these challenges, FedMosaic introduces two mechanisms, dynamic loss weighting and confidence-based aggregation.
In dynamic loss weighting, each client learns how much to trust the global signal versus its own local predictions. This tradeoff is modulated via an adaptive weight. Then, client $i$ in round $t$ minimizes the objective.
$\mathcal{L}_{i}^{t}(h)=\mathcal{L}(h,D_{i})+\lambda_{i}^{t}\cdot\mathcal{L}(h,P_{t})$
for a model $h$, where $\lambda_{i}^{t}$ is the dynamic weight and $P_{t}=(U,L_{t})$ is the pseudo-labeled public dataset in round $t$. We want $\lambda$ to be $1$ when the model performs equally well on local and global data to ensure collaboration, and to be 00 when the model performs poorly on global data, indicating that collaboration should be avoided. In cases where local data is a noisy variant of global data, $\lambda$ might even exceed $1$, but it should be bounded. To fulfill these criteria, we propose to define it as
$\lambda_{i}^{t}=\exp\left(-\frac{\mathcal{L}(h^{i}_{t-1},P_{t})-\mathcal{L}(h^{i}_{t-1},D_{i})}{\mathcal{L}(h^{i}_{t-1},D_{i})}\right)\in[0,e].$
The ratio measures the relative difference in performance on local and global data and the exponential transformation ensures that $\lambda$ is positive, smooth and bounded. This allows clients to balance collaboration and independence based on their own performance.
In confidence-based aggregation, the predictions of each client are weighed by their per-sample confidence, or expertise (in contrast to treating all clients equally when generating pseudo-labels). Clients contribute more to the global consensus when they are likely to be correct, and less when uncertain. This encourages accurate clients to lead the collaboration in regions where they are competent, while reducing the impact of noisy or misaligned peers. In practice, this means that every $b$ rounds, each client not only shares its predictions $L_{t}^{i}\in\{0,1\}^{|U|\times C}$ on the publica dataset $U$ as a one-hot encoded matrix, but also an associated confidence vector $E_{t}^{i}\in(0,\infty)^{|U|}$, where $E_{t}^{i}[j]$ quantifies the estimated reliability of client $i$’s prediction on $x_{j}$. The server collects all $(L_{t}^{i},E_{t}^{i})$ and computes a weighted score matrix
$S_{t}=\sum_{i=1}^{m}\operatorname{diag}(E_{t}^{i})\cdot L_{t}^{i}\in\mathbb{R}^{|U|\times C}.$
With this score matrix, the consensus is now formed via the confidence-based aggregation
$L_{t}[j]=\arg\max_{c\in[C]}S_{t}[j,c],\quad\forall j\in\{1,\dots,|U|\}\enspace.$
These consensus pseudo-labels are then redistributed to all clients for further training.
We propose two variants of confidence scores $E_{t}^{i}[j]$, namely a class-frequency-based heuristic, where $E_{t}^{i}[j]$ for example $x_{j}$ predicted to belong to class $y$ is the frequency of class $y$ in the local training set of client $i$, and (2) an uncertainty-based score, where $E_{t}^{i}[j]$ is inversely related to the predictive entropy or margin of $h_{t}^{i}(x_{j})$, i.e., the higher the score, the more certain the client about its prediction. Both choices allow the aggregation to selectively trust clients that are more reliable on specific regions of the data space. We evaluate both variants in Section 4. The full algorithm is provided in Algorithm 1, where differences to the original federated co-training are highlighted in green.
Input: communication period $b$, $m$ clients with local datasets $D^{1},\dots,D^{m}$ and learning algorithms $\mathcal{A}^{1},\dots,\mathcal{A}^{m}$, unlabeled public dataset $U$, total rounds $T$
Output: final models $h^{1}_{T},\dots,h^{m}_{T}$
1 initialize local models $h_{0}^{1},\dots,h_{0}^{m}$,  $P\leftarrow\emptyset$
2
3Locally at client $i$ at time $t$ do
4      compute local loss $\ell_{\mathrm{priv}}=\mathcal{L}(h^{i}_{t-1},D^{i})$
5      compute pseudo-label loss $\ell_{\mathrm{pseudo}}=\mathcal{L}(h^{i}_{t-1},P)$
6      compute adaptive weight $\lambda_{t}^{i}=\exp\left(-\frac{\ell_{\mathrm{pseudo}}-\ell_{\mathrm{priv}}}{\ell_{\mathrm{priv}}}\right)$
7      compute loss $\ell=\ell_{\mathrm{priv}}+\lambda_{t}^{i}\ell_{\mathrm{pseudo}}$
8      update $h_{t}^{i}\leftarrow\mathcal{A}^{i}(\ell,h^{i}_{t-1})$
9
10     if $t\ \%\ b=b-1$ then
11           construct prediction matrix $L_{t}^{i}\in\{0,1\}^{|U|\times C}$
12           construct expertise vector $E_{t}^{i}\in(0,\infty)^{|U|}$
13           send $(L_{t}^{i},E_{t}^{i})$ to server and receive $L_{t}$
14           $P\leftarrow(U,L_{t})$
15           end if
16
17
18
19          At server at time $t$ do
20                receive $(L_{t}^{1},E_{t}^{1}),\dots,(L_{t}^{m},E_{t}^{m})$ from clients
21                compute weighted score matrix $S_{t}=\sum_{i=1}^{m}\operatorname{diag}(E_{t}^{i})\cdot L_{t}^{i}$
22                set pseudo-labels $L_{t}[j]=\arg\max_{c\in[C]}S_{t}[j,c]\quad\text{for all }j\in\{1,\dots,|U|\}$
23                send $L_{t}$ to all clients
24
Algorithm 1 Federated Co-Training with Adaptivity and Specialization (FedMosaic)
Communication analysis:
Similar to federated co-training, in each communication round (every $b$ local steps), client $i$ sends a one-hot matrix $L_{t}^{i}\in\{0,1\}^{|U|\times C}$ to the server. In addition, it sends the confidence vector $E_{t}^{i}\in\mathbb{R}^{|U|}$. Thus, the total message size increases just by one scalar per example in $U$. The communication from server to client remains the same. Thus, the overhead in communication per round over classical federated co-training is $m|U|$, i.e., the cost of sending the confidence vectors. Similar to federated co-training, FedMosaic reduces communication over parameter sharing when $|U|$ is smaller than the number of parameters in the models, as is usually the case for deep learning (Abourayya et al., 2025).
Convergence analysis:
To provide theoretical support, we analyze the convergence behavior of FedMosaic under standard assumptions in stochastic optimization. Our goal is to characterize the rate at which each client’s objective approaches a stationary point, despite the dynamic pseudo-labeling and the heterogeneity of local objectives.
We assume standard conditions, including smoothness of the loss functions, bounded gradient variance, and bounded drift of pseudo-labels across rounds. These assumptions reflect the structure of FedMosaic, where local objectives are updated periodically but converge due to the stabilization of pseudo-labels as shown by Abourayya et al. (2025).
Assumptions 1.
The following conditions hold for each client $i\in[m]$ at round $t$:
  1. 1.
    Each loss function $\mathcal{L}_{i}^{\mathrm{local}}$ and $\mathcal{L}_{i}^{\mathrm{global},t}$ is $L(1+e)^{-1}$-smooth.
  2. 2.
    The gradient estimator $g_{i}^{t}$ is unbiased and has bounded variance:
    $\mathbb{E}[g_{i}^{t}]=\nabla\mathcal{L}_{i}^{t}(\theta_{t}),\quad\mathbb{E}[\|g_{i}^{t}-\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq\sigma^{2}.$
  3. 3.
    The global loss has bounded gradients: $\|\nabla\mathcal{L}_{i}^{\mathrm{global},t}(\theta)\|\leq G$ for all $\theta$ and $t$.
  4. 4.
    The objective drift is bounded:
    $|\mathcal{L}_{i}^{t+1}(\theta)-\mathcal{L}_{i}^{t}(\theta)|\leq\delta,\quad\forall\theta.$
  5. 5.
    The per-sample gradient variance is bounded:
    $\begin{split}\mathbb{E}_{x\in D_{i}}\left[\left\|\nabla_{\theta}\ell(\theta,x,\hat{y}^{t})-\nabla\mathcal{L}_{i}^{local,t}\right\|^{2}\right]\leq\overline{\sigma}^{2}\\ \mathbb{E}_{x\in U}\left[\left\|\nabla_{\theta}\ell(\theta,x,\hat{y}^{t})-\nabla\mathcal{L}_{i}^{global,t}\right\|^{2}\right]\leq\tilde{\sigma}^{2}\\ \end{split}$
Under these conditions, we establish that FedMosaic converges to an approximate stationary point. Specifically, after $T$ communication rounds, the average squared gradient norm decreases at a rate of $\mathcal{O}(1/T)$ plus additive terms accounting for local and global variance and pseudo-label drift.
Proposition 1 (Convergence of FedMosaic).
Let each client’s objective at round $t$ be
$\mathcal{L}_{i}^{t}(\theta)=\mathcal{L}_{i}^{\mathrm{local}}(\theta)+\lambda_{i}^{t}\mathcal{L}_{i}^{\mathrm{global},t}(\theta),\text{ where }\lambda_{i}^{t}=\exp\left(-\frac{\mathcal{L}_{i}^{\mathrm{global}}(\theta_{t})-\mathcal{L}_{i}^{\mathrm{local},t}(\theta_{t})}{\mathcal{L}_{i}^{\mathrm{local},t}(\theta_{t})}\right),$
and $\mathcal{L}_{i}^{\mathrm{global},t}$ may change at each round due to pseudo-label updates. Under Assumptions 1-5, for a fixed step size $0<\eta\leq(2L)^{-1}$ and $\min_{i}|D_{i}|=d$, after $T$ rounds of FedMosaic, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq\frac{4L\left(\mathcal{L}_{i}^{0}-\mathcal{L}_{i}^{*}\right)}{T}+\frac{\overline{\sigma}^{2}}{2Ld}+\frac{e^{2}\tilde{\sigma}^{2}}{2L|U|}+2\delta\enspace.$
The proof is provided in Appendix 6. Abourayya et al. (2025) show that under the assumption of increasing local accuracy, pseudo-labels stabilize after some round $t_{0}$, so the assumption of a bounded change in the client objective is realistic. In fact, the global loss term effectively becomes stationary under these assumptions quickly and the expected drift becomes negligibly small as $t$ increases.
Privacy considerations:
A key advantage of FedMosaic lies in its privacy-preserving design. Unlike traditional federated learning approaches that share model parameters or gradients, both known to be vulnerable to reconstruction and membership inference attacks (Zhu et al., 2019), our method shares only hard predictions on a public unlabeled dataset. This has been shown to significantly reduce privacy leakage compared to parameter- or soft-label-sharing schemes (Abourayya et al., 2025). In addition to the prediction matrix $L_{t}^{i}$, clients share a confidence vector $E_{t}^{i}$, where each element $E_{t}^{i}[j]$ quantifies their estimated expertise for the corresponding prediction. While $L_{t}^{i}$ is minimally informative, $E_{t}^{i}$ could potentially encode sensitive characteristics of the client’s private data distribution. To mitigate this risk, we propose a differentially private variant of FedMosaic (DP-FedMosaic) in which calibrated noise is added to the expertise scores before sharing. This ensures differential privacy while preserving the functional role of confidence-based collaboration. Our differentially private variant DP-FedMosaic is evaluated in sec.4. The results demonstrate that strong privacy can be guaranteed with minimal degradation in performance.
4 Empirical Evaluation
In this section, we show that FedMosaic outperforms state-of-the-art personalized federated learning methods across three challenging scenarios: (1) label skew, (2) feature shift, and (3) hybrid settings combining label and feature heterogeneity. We evaluate our method across multiple datasets and compare against strong PFL baselines, local training, and centralized upper bounds. Additional ablations and results are included in the appendix.
Label skew:
Label skew is a commonly used protocol to simulate heterogeneous output distributions, where clients observe only subsets of the label space. We consider two standard variants: (i) a pathological setting, where each client sees only a minimal subset of possible classes, and (ii) a more realistic setting where label proportions are sampled using an asymmetric distribution, inducing overlapping but imbalanced class distributions. These settings are widely adopted in the literature (T Dinh et al., 2020; Fallah et al., 2020; Zhang et al., 2023a, d, b).
In both the label skew and feature shift experiments, we use the class-frequency-based variant of the confidence score. This heuristic estimates confidence by the relative frequency of the predicted class in the local training data, which is readily available and effective when class imbalance is a primary factor of heterogeneity.
We evaluate on FashionMNIST and CIFAR10 with $m=15$ clients. In the pathological case, each client holds data from only 2 classes, resulting in extreme label imbalance. In the practical setting, label proportions are drawn from a Dirichlet distribution, simulating more natural variation. As shown in Table 2, FedMosaic consistently outperforms all baselines in both scenarios. On FashionMNIST, FedMosaic achieves $0.9940$ accuracy in the pathological setting, surpassing strong PFL methods like Ditto and APFL and even slightly outperforming both local and centralized training. On CIFAR10, FedMosaic reaches $0.8803$, again outperforming all competitors. In the practical scenario, performance trends are similar: FedMosaic achieves $0.9843$ and $0.8615$ on FashionMNIST and CIFAR10, respectively, demonstrating robustness across severity levels of label imbalance.
Table 2: Test accuracy under pathological and practical Non-IID Settings for $m=15$ clients. Color Map: baselines, worse than both baselines, worse than local training, better than both baselines.
Method Pathological non-IID Practical non-IID
FashionMNIST CIFAR10 FashionMNIST CIFAR10
Centralized $0.9928$ $(0.001)$ $0.8790$ $(0.001)$ $0.9928$ $(0.0003)$ $0.8790$ $(0.0004)$
Local training $0.9932$ $(0.0002)$ $0.8801$ $(0.0001)$ $0.9823$ $(0.0001)$ $0.8391$ $(0.002)$
FedAvg $0.7672$ $(0.001)$ $0.6442$ $(0.002)$ $0.8371$ $(0.002)$ $0.7028$ $(0.004)$
FedProx $0.7788$ $(0.003)$ $0.7025$ $(0.002)$ $0.8414$ $(0.003)$ $0.7335$ $(0.004)$
FedCT $0.7815$ $(0.0001)$ $0.7391$ $(0.0002)$ $0.8527$ $(0.0.0001)$ $0.7439$ $(0.0001)$
Per-FedAvg $0.9863$ $(0.0002)$ $0.8720$ $(0.0001)$ $0.9711$ $(0.0001)$ $0.8137$ $(0.002)$
Ditto $0.9937$ $(0.0001)$ $0.8794$ $(0.0001)$ $0.9839$ $(0.0002)$ $0.8389$ $(0.0004)$
FedBN $0.7804$ $(0.003)$ $0.8135$ $(0.005)$ $0.8539$ $(0.003)$ $0.8041$ $(0.007)$
pFedMe $0.7480$ $(0.004)$ $0.8147$ $(0.003)$ $0.8001$ $(0.001)$ $0.8161$ $(0.004)$
APFL $0.9926$ $(0.0004)$ $0.8798$ $(0.0001)$ $0.9796$ $(0.0003)$ $0.8381$ $(0.002)$
FedPHP $0.9930$ $(0.0001)$ $0.8790$ $(0.0001)$ $0.9840$ $(0.0001)$ $0.8375$ $(0.0003)$
FedMosaic $\mathbf{0.9940}$ $(0.0001)$ $\mathbf{0.8803}$ $(0.0001)$ $\mathbf{0.9843}$ $(0.0001)$ $\mathbf{0.8615}$ $(0.0001)$
Table 3: Test accuracy on Office and DomainNet datasets in feature shift scenarios: For Office-Caltech10: A, C, D ,W are abbreviations for Amazon, Caltech, DSLR and WebCam, for DomainNet: C, I, P, Q, R, S are abbreviations for Clipart, Infograph, Painting, Quickdraw, Real and Sketch. Color Map: baselines, worse than both baselines, worse than local training, better than both baselines.
Method Office DomainNet
A C D W C I P Q R S
Centralized $0.7403$ ($0.001$) $0.5824$ ($0.002$) $0.7912$ ($0.002$) $0.7852$ ($0.0001$) $0.7053$ ($0.004$) $0.3059$ ($0.003$) $0.6187$ ($0.002$) $0.7150$ ($0.001$) $0.7017$ ($0.004$) $0.6462$ ($0.003$)
Local training $0.7136$ ($0.0002$) $0.3867$ ($0.003$) $0.8125$ ($0.001$) $0.7627$ ($0.002$) $0.6531$ ($0.005$) $0.3825$ ($0.007$) $0.6652$ ($0.003$) $0.7843$ ($0.003$) $0.7104$ ($0.002$) $0.7053$ ($0.006$)
FedAvg $0.7188$ ($0.001$) $0.4844$ ($0.001$) $0.4063$ ($0.002$) $0.5424$ ($0.006$) $0.5571$ ($0.002$) $0.2842$ ($0.005$) $0.4025$ ($0.003$) $0.5264$ ($0.002$) $0.5415$ ($0.001$) $0.5612$ ($0.002$)
FedProx $0.7344$ ($0.002$) $0.5200$ ($0.002$) $0.6875$ ($0.004$) $0.7966$ ($0.004$) $0.5941$ ($0.002$) $0.3574$ ($0.004$) $0.4882$ ($0.004$) $0.5537$ ($0.001$) $0.5682$ ($0.005$) $0.5917$ ($0.002$)
FedCT $0.7396$ ($0.001$) $0.5721$ ($0.002$) $0.6873$ ($0.0001$) $0.7031$ ($0.0002$) $0.6153$ ($0.003$) $0.3519$ ($0.0001$) $0.6473$ ($0.0003$) $0.6082$ ($0.0001$) $0.7185$ ($0.0002$) $0.6925$ ($0.0001$)
Per-FedAvg $0.7304$ ($0.001$) $0.5181$ ($0.005$) $0.6922$ ($0.003$) $0.7758$ ($0.0001$) $0.6842$ ($0.0001$) $0.3621$ ($0.002$) $0.6049$ ($0.002$) $0.7263$ ($0.001$) $0.7084$ ($0.003$) $0.6816$ ($0.003$)
Ditto $0.7530$ ($0.0001$) $0.5791$ ($0.003$) $0.7839$ ($0.0002$) $0.7839$ ($0.001$) $0.7097$ ($0.0001$) $0.3913$ ($0.0001$) $0.6731$ ($0.0002$) $0.8033$ ($0.0003$) $0.7735$ ($0.0001$) $0.7314$ ($0.0003$)
FedBN $0.7539$ ($0.0001$) $0.5813$ ($0.0001$) $0.7854$ ($0.002$) $0.7823$ ($0.008$) $0.6945$ ($0.003$) $0.3801$ ($0.001$) $0.6812$ ($0.002$) $0.7921$ ($0.002$) $0.7620$ ($0.001$) $0.6923$ ($0.001$)
pFedMe $0.7083$ ($0.003$) $0.4978$ ($0.001$) $0.7500$ ($0.0003$) $0.6441$ ($0.0001$) $0.6721$ ($0.001$) $0.3742$ ($0.003$) $0.6517$ ($0.002$) $0.7524$ ($0.002$) $0.7419$ ($0.001$) $0.6893$ ($0.003$)
APFL $0.7130$ ($0.0001$) $0.3905$ ($0.0006$) $0.5085$ ($0.002$) $0.6963$ ($0.001$) $0.6873$ ($0.001$) $0.3805$ ($0.003$) $0.6739$ ($0.003$) $0.7914$ ($0.0001$) $0.7742$ ($0.001$) $0.7185$ ($0.002$)
FedPHP $0.7063$ ($0.005$) $0.4013$ ($0.0004$) $0.5178$ ($0.0001$) $0.7274$ ($0.0002$) $0.6529$ ($0.004$) $0.3632$ ($0.003$) $0.6601$ ($0.005$) $0.7703$ ($0.002$) $0.7528$ ($0.006$) $0.7011$ ($0.001$)
FedMosaic $\mathbf{0.8021}$ ($0.0001$) $\mathbf{0.6000}$ ($0.0002$) $\mathbf{0.8125}$ ($0.0002$) $\mathbf{0.8305}$ ($0.001$) $\mathbf{0.7136}$ ($0.001$) $\mathbf{0.4159}$ ($0.002$) $\mathbf{0.6938}$ ($0.004$) $\mathbf{0.8427}$ ($0.001$) $\mathbf{0.7925}$ ($0.003$) $\mathbf{0.7503}$ ($0.002$)
Feature Shift
To evaluate robustness under heterogeneous input distributions, we test FedMosaic on feature shift scenarios using Office-Caltech10 and DomainNet. In this setting, all clients share the same label space but receive data from different domains (e.g., Amazon vs. Webcam). Office-Caltech10 contains four domains (Amazon (A), Caltech(C), DSLR(D), and Webcam (W)), and DomainNet spans six domains (Clipart(C), Infograph(I), Painting(P),Quickdraw(Q),Real(R) and Sketch(S)), offering greater complexity and domain variability. While not as prominent as label skew, feature shift has scenarios have been used to evaluate FL and PFL methods Li et al. (2021c); Lee et al. (2023); Mclaughlin and Su (2024). As shown in Table 3, FedMosaic consistently achieves the highest accuracy on all domains, outperforming both traditional and personalized FL baselines. For example, on Office-Caltech10, FedMosaic attains $0.8021$ on Amazon, $0.6000$ on Caltech, $0.8125$ on DSLR, and $0.8305$ on Webcam—improving on Ditto, FedBN, and APFL. On DomainNet, FedMosaic again sets the highest scores across all six domains. These results demonstrate that our adaptive, confidence-weighted strategy effectively handles domain-specific variations while promoting generalization.
Hybrid Distribution (Label Skew + Feature Shift)
We now consider the most challenging case: combining label skew and feature shift. To simulate this hybrid heterogeneity, we partition each domain in DomainNet and Office-Caltech10 into multiple clients, assigning to each a random subset of classes. Specifically, we sample 5 clients per domain, each with access to only 2 of 10 classes (similar to the pathological label skew), resulting in 30 clients for DomainNet and 20 for Office. In this scenario, we compare both variants of our confidence-based aggregation strategy, a class-frequency-based heuristic (W) and an uncertainty-based score (U).
Table 4 shows results for multiple architectures (AlexNet and ViT). In all cases, FedMosaic significantly outperforms the best existing methods. For example, FedMosaic (U) achieves $0.8943$ accuracy on Office-Caltech10, compared to $0.8063$ for the best baseline (Ditto). On DomainNet, it reaches $0.8836$ (AlexNet) and $0.8735$ (ViT), again surpassing all alternatives. Notably, our differentially private variant (DP-FedMosaic) maintains strong performance with minimal degradation, illustrating that privacy can be ensured without compromising collaborative gains. Interestingly, both the class-frequency-based and uncertainty-based variants of FedMosaic achieve similarly strong performance across the hybrid settings. While the uncertainty-based method offers greater flexibility by capturing instance-level predictive confidence, the class-frequency heuristic performs equally well despite its simplicity. This may be due to the use of pathological label skew in our hybrid setup, where each client sees only a few classes—making frequency a strong proxy for confidence.
Taken together, these results validate that FedMosaic consistently outperforms both local and centralized baselines across diverse and realistic non-IID settings—demonstrating that adaptive, expert-aware collaboration is a principled and practical strategy for personalized federated learning.
Table 4: Test accuracy on Office and DomainNet datasets in hybrid settings scenario for $m=30$ clients on DomainNet and $m=20$ for Office dataset. Color Map: baselines, worse than both baselines, worse than local training, better than both baselines.
Method DomainNet DomainNet (ViT) Office
Centralized $0.6624$ $(0.004)$ $0.6825$ $(0.002)$ $0.4092$ $(0.006)$
Local training $0.8464$ $(0.001)$ $0.8492$ $(0.003)$ $0.8679$ $(0.004)$
FedAvg $0.3100$ $(0.008)$ $0.3328$ $(0.005)$ $0.3725$ $(0.008)$
FedProx $0.5523$ $(0.001)$ $0.5718$ $(0.003)$ $0.5839$ $(0.003)$
FedCT $0.5638$ $(0.0001)$ $0.6752$ $(0.0002)$ $0.5942$ $(0.0002)$
Per-FedAvg $0.7248$ $(0.004)$ $0.7319$ $(0.003)$ $0.7192$ $(0.005)$
Ditto $0.8147$ $(0.0001)$ $0.8382$ $(0.0002)$ $0.8063$ $(0.0001)$
FedBN $0.7154$ $(0.003)$ $0.7039$ $(0.004)$ $0.7548$ $(0.003)$
pFedMe $0.7521$ $(0.005)$ $0.7681$ $(0.008)$ $0.7483$ $(0.007)$
APFL $0.8059$ $(0.003)$ $0.8327$ $(0.005)$ $0.8091$ $(0.001)$
FedPHP $0.7825$ $(0.006)$ $0.7731$ $(0.007)$ $0.7636$ $(0.004)$
FedMosaic (W) $0.8744$ $(0.0002)$ $0.8852$ $(0.002)$ $0.8906$ $(0.0001)$
FedMosaic (U) $0.8836$ $(0.0001)$ $0.8735$ $(0.001)$ $0.8943$ $(0.0003)$
DP-FedMosaic (U) $0.8593$ $(0.001)$ $0.8574$ $(0.001)$ $0.8719$ $(0.0004)$
5 Discussion and Conclusion
Personalized Federated learning (PFL) aims to address data heterogeneity by tailoring models to client-specific distributions. Yet, as we have demonstrated, many existing approaches fall short of their promise, often failing to pass even local training or centralized baselines. This raises fundamental concerns about how collaboration is structured in personalized federated learning.
We argue that meaningful personalization in federated learning requires more than per-client modeling: it must involve adaptive, data-specific collaboration. In particular, PFL methods should support example-level decision-making, allowing clients to modulate the degree and direction of collaboration based on local context and per-sample reliability. Without this level of adaptivity, personalization risks becoming a superficial modification of global training. FedMosaic is one concrete instantiation of this principle. It enables example-level collaboration through dynamic loss weighting and confidence-based aggregation over a shared unlabeled dataset. Unlike prior methods that personalize only at the client level, FedMosaic allows each client to adapt both how much and from whom to learn, based on alignment between public and private data.
Empirical results across a diverse set of non-IID scenarios support the effectiveness of this approach. In the hybrid scenario, which combines label skew and feature shift, FedMosaic outperforms all competitors and baselines by a wide margin. In the feature shift scenarios, it again surpasses all methods across most domains, in many domains with substantial gains. In the label skew setting, FedMosaic consistently achieves the best performance for the pathological non-IID scenario, though with very narrow margins, in particular wrt. local training. In the practical non-IID scenario with milder heterogeneity, centralized training performs best, as expected. Yet, traditional federated learning methods fall short, being outperformed by several PFL approaches, including FedMosaic.
These results illustrate both strengths and limitations of personalized FL. One limitation is that, particularly in the label skew setting, the advantage over strong local baselines can be modest. Such scenarios, in particular the pathological non-IID one, raise the question of whether collaboration is justified, and whether evaluation setups with strong local baselines but weak global benefit are well-posed. We therefore emphasize the need for more meaningful benchmarks: scenarios where collaboration has clear potential upside, and where the evaluation criteria capture the practical value of federated interaction, not just statistical differences. That said, FedMosaic demonstrates that adaptive and data-aware collaboration is feasible and effective. It is the only method to outperform both local and centralized baselines across all settings, validating its robustness and practical utility.
While FedMosaic represents a principled and practically validated advance in personalized federated learning, it also opens new directions for future work. A key limitation is the assumption of a public unlabeled dataset. Although such datasets exist in many domains, e.g., healthcare, vision, and language, it remains an open question how to extend this paradigm when such data are limited or unavailable. Developing mechanisms for privacy-preserving dataset synthesis, or leveraging foundation models for public data distillation, could further broaden the applicability of our framework.
While we provide a differentially private variant of FedMosaic, a deeper theoretical understanding of the trade-offs between privacy, collaboration, and generalization remains an open challenge. These questions go beyond our method, they point toward a more general agenda of making federated collaboration both trustworthy and adaptive. We see this work as the starting point of a broader research agenda: developing federated learning algorithms that adaptively personalize collaboration itself, not just models, and doing so in a way that is data-aware, privacy-preserving, and robust across deployment conditions.
A Note on the Byzantine Resilience of FedMosaic
Following the argument by [Jiang et al., 2020], who show that federated semi-supervised learning with soft labels sharing (e.g., FedDistill) is more Byzantine resilient than FedAvg due to the bounded nature of the threat vector on the probability simplex, we argue that FedMosaic exhibits similar ( if not stronger) resilience properties. Like FedCT [Abourayya et al., 2025], FedMosaic relies on hard label sharing, further constraining the threat vector to a binary classification decision per example. Moreover, FedMosaic incorporates confidence-based aggregation, which naturally downweights unreliable predictions. This mechanism provides an additional layer of robustness by reducing the influence of low confidence ( and potentially malicious) clients. While a formal analysis remains open, these properties suggest that FedMosaic may be at least as Byzantine resilient as FedDistill and FedCT. Exploring this direction further is promising for future work.
8 Details on Experiments
Label Skew
FashionMinst and CIFAR10 datasets have been used for label skew experiments. In FashionMinst, we converted the raw grayscale $28\times 28$ images into Pytorch tensors and normalized pixel values to the range $[-1,1]$ using a mean of $0.5$ and standard deviation of $0.5$. In CIFAR10, we converted RGB $32\times 32$ images into Pytorch tensors of shape $[3,32,32]$ and normalizes each color channel independently to the range of $[-1,1]$, using a mean of $0.5$ and standard deviation of $0.5$. The data is partitioned across $15$ clients. In a pathological non-IID setting, each client receives data from only $2$ out of $10$ classes. In a practical non-IID setting, data is distributed across $15$ clients using a Dirichlet distribution. This creates naturally overlapping, imbalanced label distributions among clients. Training data distribution of each scenario of CIFAR10 are showing in Fig.7 and Fig.7.
Figure 6: CIFAR10 clients data distribution
in Pathological non-IID setting
Figure 7: CIFAR10 clients data distribution
in Practical non-IID setting
Feature Shift
we used the Office-Caltech10 and DominNet datasets. For both, we adopt AlexNet as a neural network architecture. Input images are resized to $256\times 256\times 3$. Training is performed till convergence using the corss-entropy loss and Adam optimizer with learning rate of $10^{-2}$. We use a batch size of $32$ for Office-Caltech10 dataset and $64$ for DomainNet. For DomainNet, which originally contains $345$ categories, we restrict the label space to the top $10$ most frequent classes to reduce complexity, The selected categories are: bird, feather, headphones, icecream, teapot, tiger, whale, windmill, wineglass, zebra. For Office-Caltech10, each client get one of the $4$ domains and For DomainNet dataset, each client get one of the $6$ domains. The distribution of each client training data are showing in Fig.9 and Fig.9.
Figure 8: DomainNet clients data distribution.
Figure 9: Office-Caltech10 clients data distribution.
Hybrid Distribution
We simulate the hybrid data distribution by combining both label distribution skew and feature distribution shift. We use the same two datasets as in feature shift experiments: Office-Caltech10 and DomainNet. To introduce label skew, for each domain, we randomly sample $5$ clients and assign to each client only $2$ out of $10$ total classes. This results in $20$ clients for the Office-Caltech10 dataset ($4$ domains $\times$ $5$ clients) and $30$ clients for DomainNet ($6$ domains $\times$ $5$ clients). This creates a hybrid non-IID setting where clients differ significantly in both input distribution and output distribution. We use the same preprocessing and training configurations as the feature shift experiments. All input images are resized to $256\times 256\times 3$ before being fed into $AlexNet$. Models are trained using cross-entropy loss and Adam optimizer with learning rate of $10^{-2}$. The batch size is set to $32$ for Office-Caltech10 and $64$ for DomainNet. For DomainNet, we selected the $10$ most frequent as feature shift experiments.To effectively visualize the distribution of local training data across $30$ clients, we used a dot matrix plot, which offers a compact and intuitive representation of client-level variation. The visualization of the Clients distribution of DomainNet and Office-Caltech10 datasets are shown in Fig.11 and Fig.11
Figure 10: DomainNet clients Hybrid data distribution.
Figure 11: Office-Caltech10 clients Hybrid data distribution.
9 Practical Impact of FedMosaic
FedMosaic demonstrates substantial practical impacts in personalized federated learning (PFL), specifically addressing data heterogeneity across decentralized clients. FedMosaic uses a fine-grained collaboration mechanism that enables each client to selectively rely on the collective expertise available, enhancing model accuracy and robustness. This approach is particularly valuable in domains with significant data variability, such as healthcare, finance, and personalized recommendation systems, where traditional federated methods often fail. Empirical results highlight that FedMosaic consistently outperforms state-of-the-art PFL methods, as well as local and centralized baselines across various challenging scenarios, including label skew, feature shift, and hybrid heterogeneity distribution. Its design inherently preserves privacy by limiting data exchanges to hard predictions over a shared unlabeled dataset, thus minimizing potential privacy leakage compared to parameter sharing approaches. This follows the approach to share as little as possible [Mian et al., 2023, Tan et al., 2022] and therefore adheres to the privacy-by-design-principle [Cavoukian et al., 2009]. Furthermore, FedMosaic’s differentially private variant (DP-FedMosaic) showcases the feasibility of achieving robust privacy guarantees without significant compromise on performance, making it particularly attractive for sensitive applications. Another aspect is the communication efficiency of federated co-training approaches for large models: in those cases the number of parameters vastly exceed the number of labels for the public dataset. Combining this with communication-efficient protocols [Kamp et al., 2016, Kamp, 2019] has the potential to reduce communication by several orders of magnitude, in particular for large transformer-based models, such as LLMs.